### v2

@@
Distributionally Robust Dynamic Structural Estimation: Serial Dependence and Sensitivity Analysisthanks: I would like to thank Lars Nesheim, Dennis Kristensen, and Aureo de Paula for their continuous guidance and support. I am grateful to Victor Aguirregabiria, Dante Amengual, Debopam Bhattacharya, Stephane Bonhomme, Christian Bontemps, Cuicui Chen, Tim Christensen, Kevin Dano, Ben Deaner, Ivan Fernandez-Val, Hugo Freeman, Alfred Galichon, Raffaella Giacomini, Gautam Gowrisankaran, Joao Granja, Jiaying Gu, Sukjin Han, Stefan Hubner, Koen Jochmans, Hiroaki Kaido, Hiro Kasahara, Yuichi Kitamura, Adam Lee, Arthur Lewbel, Lorenzo Magnolfi, Chuck Manski, Matt Masten, Robert McCann, Adam McCloskey, Nour Meddahi, Bob Miller, Marcel Nutz, Ariel Pakes, Silvana Pesenti, Lei Qiao, Mathias Reynaert, Eduardo Souza-Rodrigues, Pietro Emilio Spini, Jörg Stoye, Elie Tamer, Xun Tang, Weining Wang, Martin Weidner, Leonard Wong, Andrei Zeleneev for helpful discussions and comments. I also thank seminar participants at UCL, TSE, and University of Toronto, as well as conference audiences at 2025 Bristol Econometric Study Group, 2025 Midwest Econometrics Group Conference, and IAAE 2025.
@@
(November 12, 2025)
Distributional assumptions that discipline serially correlated latent variables play a central role in dynamic structural models. We propose a framework to quantify the sensitivity of scalar parameters of interest (e.g., welfare, elasticity) to such distributional assumptions. We derive bounds on the scalar parameter by perturbing a reference distribution, while imposing a stationarity condition for time-homogeneous models or a Markovian condition for time-inhomogeneous models. The bounds are the solutions to optimization problems, for which we derive a computationally tractable dual formulation. We establish consistency, convergence rate, and asymptotic distribution for the estimator of the bounds. We demonstrate the approach with two applications: an infinite-horizon dynamic demand model for new cars in the United Kingdom, Germany, and France, and a finite-horizon dynamic labor supply model for taxi drivers in New York City. In the car application, perturbed price elasticities deviate by at most 15.24% from the reference elasticities, while perturbed estimates of consumer surplus from an additional \$3,000 electric vehicle subsidy vary by up to 102.75%. In the labor supply application, the perturbed Frisch labor supply elasticity deviates by at most 76.83% for weekday drivers and 42.84% for weekend drivers.
Keywords: Serial dependence, Sensitivity analysis, Dynamic structural models, Robustness, Misspecification
@@ 1 Introduction
Dynamic structural models are useful tools for counterfactual policy analysis in various fields of economics. The correlation structure of latent variables is a key feature of these models, as it captures the persistence of unobserved factors that affect agents’ decisions over time. Examples of potentially serially correlated latent variables include product characteristics in demand estimation (nair2007intertemporal; schiraldi2011automobile; gowrisankaran2012dynamics), search costs in consumer search (koulayev2014search), firm productivity in trade (piveteau2021empirical), patent profitability in optimal stopping (pakes1984patents), quality in technology adoption (de2019subsidies), health shocks in insurance (fang2021life), and beliefs about ability in labor economics (miller1984job; arcidiacono2025college).
@@
Distributional assumptions about the serial dependence of latent variables play a central role in the estimation and counterfactual analysis of dynamic structural models. These assumptions capture agents’ uncertainty about the future, such as a consumer’s uncertainty about future product characteristics in demand estimation. The misspecification of these assumptions can lead to biased estimates of future continuation value, which in turn bias counterfactual predictions (e.g., welfare and elasticity). However, the direction and magnitude of this bias are unclear, because the models are dynamic and nonlinear. This raises the need for sensitivity analysis of empirical results to these distributional assumptions.
@@
In this paper, we propose a framework to quantify this sensitivity by perturbing a reference distribution to compute bounds on a scalar parameter of interest. The scalar parameter (e.g., welfare and elasticity) is a function of model primitives, such as model parameters, the distribution of latent variables, and the value function in dynamic structural models. In our proposed framework, the bounds on this scalar parameter are the solutions to constrained optimization problems whose feasible region (identified set) is defined by moment conditions for estimation, structural constraints (e.g., the Bellman equation), and the perturbation set around the reference distribution. The distribution that attains the bound is called the worst-case distribution.
@@
A central challenge is to define the perturbation set in a way that simplifies computation while maintaining key structural features of the distribution of latent variables. For time-homogeneous models, the structural feature is the stationarity condition, i.e., the perturbed transition distribution must be stationary. For finite-horizon, time-inhomogeneous models, the perturbed trajectory of latent variables must be Markovian. In addition, the terminal distribution is fixed because it can be nonparametrically point identified (lewbel2000semiparametric; matzkin2007nonparametric). Because the stationarity and Markov conditions are functional constraints imposed directly on the distribution being optimized, they are computationally difficult to impose. We contribute to the sensitivity analysis literature (e.g., christensen2023counterfactual) by providing a computationally tractable framework to deal with these constraints.
@@
To address these challenges, we define the perturbation set as a Kullback-Leibler (KL) divergence ball around the reference distribution. The KL radius controls the size of the perturbation set. Then, we employ the Optimal Transport (OT) framework to impose constraints on the distribution of latent variables. These constraints, combined with the KL divergence constraint, form a computationally tractable entropic optimal transport (EOT) problem. For the time-homogeneous case, the reference distribution is a joint distribution of current and future latent variables. The stationarity condition requires that the marginal distributions of current and future latent variables of the perturbed joint distribution are the same, which can be imposed using OT. This marginal constraint, when combined with the KL divergence penalty from the problem’s Lagrangian, forms an EOT problem. For the time-inhomogeneous case, the perturbed trajectory must be Markovian, and the terminal distribution is fixed. We show how to formulate this as an EOT problem.
@@
Then, we consider three complementary sensitivity measures to interpret the results. First, the global sensitivity approach computes the largest deviation from the reference value. It progressively increases the KL radius until the bounds flatten. We show that it provides a tractable approximation to the nonparametric bounds on the scalar parameter when the KL divergence constraint is removed. Moreover, we derive an explicit upper bound on the approximation error, which enables us to control the error within a desired level. Second, the local sensitivity approach analyzes the effect of small perturbations. It computes the right derivatives of the bounds with respect to the KL radius at zero, which is our local sensitivity measure. Finally, the robustness metric approach, inspired by spini2021robustness, computes the smallest deviation from the reference distribution required to produce sensitive results. It is the smallest KL divergence from the reference distribution required for the scalar parameter’s value to fall below a user-specified threshold (e.g., 5% below the reference value). In addition to our three sensitivity measures, we can also estimate an alternative model and set the radius as the KL divergence between the alternative and reference models.
@@
We apply our framework to an infinite-horizon dynamic demand model for new cars in the UK, Germany, and France. We consider the sensitivity of the price elasticity and consumer surplus from an additional \$3,000 electric vehicle subsidy. In the model, the indirect utility of purchasing is a latent variable due to unobserved product characteristics, and its transition is typically modeled as an AR(1) process (e.g., schiraldi2011automobile; gowrisankaran2012dynamics). For the price elasticity, we find that the French market is the least sensitive to the distributional assumption (at most 6.20% deviation from the reference elasticity), while the German market is the most sensitive (at most 15.24% deviation). We also find that this sensitivity is relatively stable over time for all three markets. For the consumer surplus, the German market is also the most sensitive (at most 102.75% deviation from the reference consumer surplus), while the UK and French markets are less sensitive (at most 25.17% and 24.73% deviation, respectively). More importantly, the results are still economically meaningful even under the worst-case distribution, with the consumer surplus remaining at least \$309 million for the German market, \$1,243 million for the French market, and \$2,584 million for the UK market.
We also apply our framework to a finite-horizon dynamic labor supply model for taxi drivers in New York City. We consider the sensitivity of the elasticity of stopping work and the Frisch elasticity of labor supply. In the model, the market-level supply shock is a latent variable. The reference distribution is also modeled as an AR(1) process. For the elasticity of stopping work, we find that weekday drivers’ elasticity is more sensitive in the morning, while weekend drivers’ elasticity is more sensitive in the afternoon. For the Frisch elasticity, both weekday and weekend drivers’ elasticities are sensitive to the distributional assumption, with at most 76.83% and 42.84% deviation from the reference elasticity, respectively.
@@ 1.1 Related Literature
Our framework contributes to this literature in three ways. First, we focus on a scalar parameter (outcome of interest) directly instead of the identified set of model primitives. Second, to analyze sensitivity, we consider a perturbation set around the reference distribution rather than the set of all distributions. Third, we complement identification strategies that do not rely on distributional assumptions. In the labor supply application, we use the ECCP estimator by kalouptsidi2021linear to point identify the utility parameters, and then conduct sensitivity analysis of the labor supply elasticity to the distributional assumption about the market-level supply shock.
@@
Sensitivity analysis and robustness. In DDC models, kalouptsidi2021counterfactual; kalouptsidi2021identification relax common normalization on the utility function. bugni2019inference considers the local misspecification of the transition density of observable variables, assuming the transition density is correctly specified in the limit (as the sample size goes to infinity) in DDC models. andrews2017measuring considers a setting in which the moments are locally misspecified under the reference distribution. Subsequent work, armstrong2021sensitivity, constructs near-optimal confidence intervals in such models. kitamura2013robustness considers the robust estimation under moment restrictions. bonhomme2022minimizing perturbs the reference model and assumes the size of the perturbation shrinks to zero as the sample size goes to infinity. chen2024robust relaxes the rational expectation assumption. gu2023dual considers the identification of scalar counterfactual parameters using optimal transport. spini2021robustness studies the robustness of policy effects to changes in the distribution of covariates. armstrong2025misspecification provides a selective review of misspecification in econometrics. Most closely related is christensen2023counterfactual, who conducts sensitivity analysis with respect to parametric assumptions about the distribution of latent variables in structural models. However, they focus on relaxing the marginal distribution assumption while maintaining serial independence.
@@
We contribute to this literature in three ways. First, our focus on serial dependence complements analysis of misspecification in the marginal distribution assumption (e.g., christensen2023counterfactual). Second, the size of our perturbation set is fixed and does not shrink with the sample size. Third, we provide a computationally tractable dual formulation for the optimization problem. While prior work (schennach2014entropic; gu2023dual; christensen2023counterfactual) uses duality to convert the infinite-dimensional problem to a finite-dimensional one, our dual problem is still infinite-dimensional. This is because the stationarity condition for time-homogeneous models and fixed terminal distribution for time-inhomogeneous models are infinite-dimensional, and serial dependence leads to an infinite-dimensional value function in dynamic structural models. However, by leveraging EOT duality, we provide a tractable implementation and demonstrate our approach through two empirical applications.
@@
Our work is distinct in that we study the misspecification due to assumptions about the serial dependence of latent variables—a problem of model specification rather than data limitation. In addition, because our framework can be treated as a moment-constrained DRO problem, we employ the minimax theorem (see fan1953minimax and ricceri1998minimax, Theorem 1.3) to exchange the order of the supremum over Lagrangian multipliers (for the moment conditions and structural constraints) and the infimum over distributions in the perturbation set. This exchange allows a direct application of duality results from the DRO literature, which simplifies our proof of the dual formulation significantly.
@@
Notation: Let $U\in\mathcal{U}\subseteq\mathbb{R}^{d}$ be a vector of latent variables with support $\mathcal{U}$ where $\mathcal{U}$ is assumed to be Polish. Let $\mathcal{P}(\mathcal{U})$ be the space of Borel probability measures on $\mathcal{U}$ and $\mathcal{B}(\mathcal{U})$ be the Borel $\sigma$-algebra on $\mathcal{U}$. In this paper, all measures are assumed to be absolutely continuous with respect to the Lebesgue measure. Denote by $\mathbb{E}_{F}\left[\cdot\right]$, $\mathbb{E}_{x}\left[\cdot\right]$ the expectations with respect to $F\in\mathcal{P}(\mathcal{U})$ and the probability distribution of the random variable $x$, respectively. For variables in a stationary dynamic context, a prime (e.g., $x^{\prime}$) denotes the variable’s value in the next period. Let $F_{1},F_{2}\in\mathcal{P}(\mathcal{U})$, we write $F_{1}\ll F_{2}$ if $F_{1}$ is absolutely continuous with respect to $F_{2}$, and $F_{1}\otimes F_{2}$ as the product measure. For a finite-dimensional vector, denote by $\|\cdot\|_{p}$ the $p$-norm. Denote by $L^{p}(F)$ the space of functions for which $\int|f|^{p}dF<\infty$. For a set $\mathcal{A}$, denote by $\text{int}(\mathcal{A})$ its interior. Denote by $\mathbb{R}_{+}$ the set of non-negative real numbers, and $\mathbb{N}$ the set of natural numbers. Finally, let $\mathcal{J}:=\left\{1,\cdots,J\right\}$.

### v3

@@
Robust Structural Estimation under Misspecified Latent-State Dynamicsthanks: I would like to thank Lars Nesheim, Dennis Kristensen, and Aureo de Paula for their continuous guidance and support. I am grateful to Victor Aguirregabiria, Dante Amengual, Debopam Bhattacharya, Stephane Bonhomme, Christian Bontemps, Cuicui Chen, Tim Christensen, Kevin Dano, Ben Deaner, Ivan Fernandez-Val, Hugo Freeman, Alfred Galichon, Raffaella Giacomini, Gautam Gowrisankaran, Joao Granja, Jiaying Gu, Sukjin Han, Stefan Hubner, Koen Jochmans, Hiroaki Kaido, Hiro Kasahara, Yuichi Kitamura, Adam Lee, Arthur Lewbel, Lorenzo Magnolfi, Chuck Manski, Matt Masten, Robert McCann, Adam McCloskey, Nour Meddahi, Bob Miller, Marcel Nutz, Ariel Pakes, Silvana Pesenti, Lei Qiao, Mathias Reynaert, Eduardo Souza-Rodrigues, Pietro Emilio Spini, Jörg Stoye, Elie Tamer, Xun Tang, Weining Wang, Martin Weidner, Leonard Wong, Andrei Zeleneev for helpful discussions and comments. I also thank seminar participants at UCL, TSE, and University of Toronto, as well as conference audiences at 2025 Bristol Econometric Study Group, 2025 Midwest Econometrics Group Conference, and IAAE 2025.
@@
(November 13, 2025)
Estimation and counterfactual analysis in dynamic structural models rely on assumptions about the dynamic process of latent variables, which may be misspecified. We propose a framework to quantify the sensitivity of scalar parameters of interest (e.g., welfare, elasticity) to such assumptions. We derive bounds on the scalar parameter by perturbing a reference dynamic process, while imposing a stationarity condition for time-homogeneous models or a Markovian condition for time-inhomogeneous models. The bounds are the solutions to optimization problems, for which we derive a computationally tractable dual formulation. We establish consistency, convergence rate, and asymptotic distribution for the estimator of the bounds. We demonstrate the approach with two applications: an infinite-horizon dynamic demand model for new cars in the United Kingdom, Germany, and France, and a finite-horizon dynamic labor supply model for taxi drivers in New York City. In the car application, perturbed price elasticities deviate by at most 15.24% from the reference elasticities, while perturbed estimates of consumer surplus from an additional \$3,000 electric vehicle subsidy vary by up to 102.75%. In the labor supply application, the perturbed Frisch labor supply elasticity deviates by at most 76.83% for weekday drivers and 42.84% for weekend drivers.
Keywords: Dynamic structural models, Sensitivity analysis, Misspecified dynamics
@@ 1 Introduction
Dynamic structural models are useful tools for counterfactual policy analysis in various fields of economics. The dynamic process of latent variables is a key feature of these models, as it captures the persistence of unobserved factors that affect agents’ decisions over time. Examples of potentially serially correlated latent variables include product characteristics in demand estimation (nair2007intertemporal; schiraldi2011automobile; gowrisankaran2012dynamics), search costs in consumer search (koulayev2014search), firm productivity in trade (piveteau2021empirical), patent profitability in optimal stopping (pakes1984patents), quality in technology adoption (de2019subsidies), health shocks in insurance (fang2021life), and beliefs about ability in labor economics (miller1984job; arcidiacono2025college).
@@
Assumptions about the dynamic process governing the serial dependence of latent variables are central to the estimation and counterfactual analysis in dynamic structural models. These assumptions capture agents’ uncertainty about the future, such as a consumer’s uncertainty about future product characteristics in demand estimation. The misspecification of these assumptions can lead to biased estimates of future continuation value, which in turn bias counterfactual predictions (e.g., welfare and elasticity). However, the direction and magnitude of this bias are unclear, because the models are dynamic and nonlinear. This raises the need for sensitivity analysis of empirical results to these distributional assumptions.
@@
In this paper, we propose a framework to quantify this sensitivity by perturbing a reference dynamic process to compute bounds on a scalar parameter of interest. The scalar parameter (e.g., welfare and elasticity) is a function of model primitives, such as model parameters, the distribution of latent variables, and the value function in dynamic structural models. In our proposed framework, the bounds on this scalar parameter are the solutions to constrained optimization problems whose feasible region (identified set) is defined by moment conditions for estimation, structural constraints (e.g., the Bellman equation), and the perturbation set around the reference transition distribution. The distribution that achieves the bound is called the worst-case distribution.
@@
A central challenge is to define the perturbation set in a way that simplifies computation while maintaining key structural features of the distribution of latent variables. For time-homogeneous models, the structural feature is the stationarity condition, i.e., the perturbed dynamic process must be stationary. For finite-horizon, time-inhomogeneous models, the perturbed trajectory of latent variables must be Markovian. In addition, the terminal distribution is fixed because it can be nonparametrically point identified (lewbel2000semiparametric; matzkin2007nonparametric). Because the stationarity and Markov conditions are functional constraints imposed directly on the distribution being optimized, they are computationally difficult to impose. We contribute to the sensitivity analysis literature (e.g., christensen2023counterfactual) by providing a computationally tractable framework to deal with these constraints.
@@
To address these challenges, we define the perturbation set as a Kullback–Leibler (KL) divergence ball around the reference distribution. For the time-homogeneous case, the reference distribution is a joint distribution of current and future latent variables. For the time-inhomogeneous case, it is the joint distribution of the entire trajectory of latent variables. The KL radius controls the size of the perturbation set. We then employ the Optimal Transport (OT) framework to impose structural constraints on the distribution of latent variables. In the time-homogeneous case, the stationarity condition requires that the marginal distributions of current and future latent variables of the perturbed joint distribution coincide, which can be imposed using OT. This marginal constraint, together with the KL divergence penalty from the problem’s Lagrangian, yields a computationally tractable Entropic Optimal Transport (EOT) problem. In the time-inhomogeneous case, the perturbed trajectory must be Markovian, and the terminal distribution is fixed. We show how to formulate this as an EOT problem.
@@
Then we consider three complementary sensitivity measures to interpret the results. First, the global sensitivity approach computes the largest deviation from the reference value. It progressively increases the KL radius until the bounds flatten. We show that it provides a tractable approximation to the nonparametric bounds on the scalar parameter when the KL divergence constraint is removed. Moreover, we derive an explicit upper bound on the approximation error, which enables us to control the error within a desired level. Second, the local sensitivity approach analyzes the effect of small perturbations. It computes the right derivatives of the bounds with respect to the KL radius at zero, which serves as our local sensitivity measure. Finally, the robustness metric approach, inspired by spini2021robustness, computes the smallest deviation from the reference distribution required to produce sensitive results. It is the smallest KL divergence from the reference distribution required for the scalar parameter’s value to fall below a user-specified threshold (e.g., 5% below the reference value). In addition to our three sensitivity measures, we can also estimate an alternative model and set the radius as the KL divergence between the alternative and reference models.
@@
We apply our framework to an infinite-horizon dynamic demand model for new cars in the UK, Germany, and France. We consider the sensitivity of the price elasticity and consumer surplus from an additional \$3,000 electric vehicle subsidy. In the model, the indirect utility of purchasing is a latent variable due to unobserved product characteristics, and its transition is typically modeled as an AR(1) process (e.g., schiraldi2011automobile; gowrisankaran2012dynamics). For the price elasticity, we find that the French market is the least sensitive to the distributional assumption (at most 6.20% deviation from the reference elasticity), while the German market is the most sensitive (at most 15.24% deviation). We also find that this sensitivity is relatively stable over time for all three markets. For the consumer surplus, the German market is also the most sensitive (at most 102.75% deviation from the reference consumer surplus), while the UK and French markets are less sensitive (at most 25.17% and 24.73% deviation, respectively). Importantly, the results remain economically meaningful even under the worst-case distribution, with the consumer surplus remaining at least \$309 million for the German market, \$1,243 million for the French market, and \$2,584 million for the UK market.
We also apply our framework to a finite-horizon dynamic labor supply model for taxi drivers in New York City. We consider the sensitivity of the elasticity of stopping work and the Frisch elasticity of labor supply. In the model, the market-level supply shock is a latent variable, and its transition is also modeled as an AR(1) process. For the elasticity of stopping work, we find that weekday drivers’ elasticity is more sensitive in the morning, while weekend drivers’ elasticity is more sensitive in the afternoon. For the Frisch elasticity, both weekday and weekend drivers’ elasticities are sensitive to the distributional assumption, with at most 76.83% and 42.84% deviation from the reference elasticity, respectively.
@@ 1.1 Related Literature
Our framework contributes to this literature in three ways. First, we focus directly on a scalar parameter (e.g., welfare, elasticity) rather than on the identified set of model primitives. Second, to analyze sensitivity, we consider a perturbation set around the reference distribution rather than the set of all distributions. Third, we complement identification strategies that do not rely on distributional assumptions. In the labor supply application, we use the ECCP estimator of kalouptsidi2021linear to point identify the utility parameters and then conduct sensitivity analysis of the labor supply elasticity with respect to assumptions about the dynamic process of the market-level supply shock.
@@
Sensitivity analysis and robustness. In DDC models, kalouptsidi2021counterfactual; kalouptsidi2021identification relax common normalizations on the utility function. bugni2019inference considers the local misspecification of the transition density of observable variables, assuming the transition density is correctly specified in the limit (as the sample size goes to infinity) in DDC models. andrews2017measuring considers a setting in which the moments are locally misspecified under the reference distribution. Subsequent work armstrong2021sensitivity constructs near-optimal confidence intervals in such models. kitamura2013robustness considers the robust estimation under moment restrictions. bonhomme2022minimizing perturbs the reference model and assumes the size of the perturbation shrinks to zero as the sample size goes to infinity. chen2024robust relaxes the rational expectation assumption. gu2023dual considers the identification of scalar counterfactual parameters using optimal transport. spini2021robustness studies the robustness of policy effects to changes in the distribution of covariates. armstrong2025misspecification provides a selective review of misspecification in econometrics. Most closely related is christensen2023counterfactual, who conducts sensitivity analysis with respect to parametric assumptions about the distribution of latent variables in structural models. However, they focus on relaxing the marginal distribution assumption while maintaining serial independence.
@@
We contribute to this literature in three ways. First, our focus on misspecified dynamic processes complements previous analysis of misspecification in dynamic structural estimation (e.g., bugni2019inference; kalouptsidi2021counterfactual; kalouptsidi2021identification; christensen2023counterfactual). Second, the size of our perturbation set is fixed and does not shrink with the sample size. Third, we provide a computationally tractable dual formulation for the optimization problem. While prior work (schennach2014entropic; gu2023dual; christensen2023counterfactual) uses duality to convert the infinite-dimensional problem to a finite-dimensional one, our dual problem is still infinite-dimensional. This is because the stationarity condition for time-homogeneous models and fixed terminal distribution for time-inhomogeneous models are infinite-dimensional, and serial dependence leads to an infinite-dimensional value function in dynamic structural models. However, by leveraging EOT duality, we provide a tractable implementation and demonstrate our approach through two empirical applications.
@@
Our work is distinct in that we study the misspecification due to assumptions about the serial dependence of latent variables—a problem of model specification rather than data limitation. Moreover, because our framework can be treated as a moment-constrained DRO problem, we employ the minimax theorem (see fan1953minimax and ricceri1998minimax, Theorem 1.3) to exchange the order of the supremum over Lagrangian multipliers (for the moment conditions and structural constraints) and the infimum over distributions in the perturbation set. This exchange allows a direct application of duality results from the DRO literature, which simplifies our proof of the dual formulation significantly.
@@
Notation: Let $U\in\mathcal{U}\subseteq\mathbb{R}^{d}$ be a vector of latent variables with support $\mathcal{U}$ where $\mathcal{U}$ is assumed to be Polish. Let $\mathcal{P}(\mathcal{U})$ be the space of Borel probability measures on $\mathcal{U}$ and $\mathcal{B}(\mathcal{U})$ be the Borel $\sigma$-algebra on $\mathcal{U}$. In this paper, all measures are assumed to be absolutely continuous with respect to the Lebesgue measure. Denote by $\mathbb{E}_{F}\left[\cdot\right]$, $\mathbb{E}_{x}\left[\cdot\right]$ the expectations with respect to $F\in\mathcal{P}(\mathcal{U})$ and the probability distribution of the random variable $x$, respectively. For variables in a stationary dynamic context, a prime (e.g., $x^{\prime}$) denotes the variable’s value in the next period. Let $F_{1},F_{2}\in\mathcal{P}(\mathcal{U})$, we write $F_{1}\ll F_{2}$ if $F_{1}$ is absolutely continuous with respect to $F_{2}$, and $F_{1}\otimes F_{2}$ as the product measure. For a finite-dimensional vector, denote by $\|\cdot\|_{p}$ the $p$-norm. Denote by $L^{p}(F)$ the space of functions for which $\int|f|^{p}dF<\infty$. For a set $\mathcal{A}$, let $\text{int}(\mathcal{A})$ denote its interior. Denote by $\mathbb{R}_{+}$ the set of non-negative real numbers, and $\mathbb{N}$ the set of natural numbers. Finally, let $\mathcal{J}:=\left\{1,\cdots,J\right\}$.